# Singular dynamics for discrete weak K.A.M. solutions of exact twist maps

### v1

@@ 3.2. $\Sigma_{+}$ is Lipschitz
Take $x\in\mathbb{R}$ and denote $\Sigma_{+}(x)$ by $y$. If $(\Sigma_{+})^{-1}(y)$ is a singleton $x$, then take $x^{\prime}=x^{\prime\prime}=x$. If $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed interval $[x_{-},x_{+}]$, then take $x^{\prime}=x_{-}$ and $x^{\prime\prime}=x_{+}$. ∎
@@
Let $u$ be a discrete weak K.A.M. solution. If $x\in\pi_{1}(\mathcal{NS}(u))$, then
  1. (i)
    $u(x)-\bar{S}=T^{+}[u](x)$;
  2. (ii)
    Both $u$ and $T^{+}[u]$ are differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)$.
Proof.
Let $x\in\pi_{1}(\mathcal{NS}(u))$, then there exists $y\in\mathbb{R}$ such that $(x,y)\in\mathcal{NS}(u)$. Hence $u(x)-\bar{S}=T^{+}[u](x)$ by Remark 2.11. Pick $w\in(\Sigma_{+})^{-1}(x)$, i.e., $T^{+}[u](w)=u(x)-S(w,x)$. Combine these two equations and we get
$T^{+}[u](x)-T^{+}[u](w)=\left(u(x)-\bar{S}\right)-\left(u(x)-S(w,x)\right)=S(w,x)-\bar{S}.$
That is, $(w,x)\in\mathcal{NS}(T^{+}[u])$. Hence by (b) of Lemma 2.12, we have both $u$ and $T^{+}[u]$ are differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)$. ∎
Lemma 3.8.
If $u$ is a discrete weak K.A.M. solution, then $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u$ is Lipschitz on $\pi_{1}(\mathcal{NS}(u))$. That is, there exists $K>0$ such that
$|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})|\leq K|x_{1}-x_{2}|,\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).$
Proof.
Since $u$ is a semiconcave function, by (i) of Proposition 2.2, there exists a constant $C\geq 0$ such that
$u(x_{1})-u(x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})(x_{1}-x_{2})\leq\frac{C}{2}|x_{1}-x_{2}|^{2}$
for any $x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u))$. On the other hand, since $T^{+}[u]$ is semiconvex, there exists $C^{\prime}\geq 0$ such that
$T^{+}[u](x_{1})-T^{+}[u](x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x_{2})(x_{1}-x_{2})\geq-\frac{C^{\prime}}{2}|x_{1}-x_{2}|^{2}$
for any $x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u))$. Using Lemma 3.7 to replace $T^{+}[u]$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u]$ with $u$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u$ respectively, we have
$u(x_{1})-u(x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})(x_{1}-x_{2})\geq-\frac{C^{\prime}}{2}|x_{1}-x_{2}|^{2}$
for any $x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u))$. Therefore we have
$-\frac{C^{\prime}}{2}|x_{1}-x_{2}|^{2}\leq u(x_{1})-u(x_{2})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})(x_{1}-x_{2})\leq\frac{C}{2}|x_{1}-x_{2}|^{2},\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).$
Exchange the roles of $x_{1}$ and $x_{2}$, then we get
$-\frac{C^{\prime}}{2}|x_{1}-x_{2}|^{2}\leq u(x_{2})-u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})(x_{2}-x_{1})\leq\frac{C}{2}|x_{1}-x_{2}|^{2},\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).$
Add these two inequalities together, we get
$-C^{\prime}|x_{1}-x_{2}|^{2}\leq(\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2}))(x_{1}-x_{2})\leq C|x_{1}-x_{2}|^{2},\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).$
Hence if $K=\max\{C,C^{\prime}\}$, then
$|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2})|\leq K|x_{1}-x_{2}|,\quad\forall x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u)).\qquad\qquad\qed$
Proposition 3.9.
If $u$ is a discrete weak K.A.M. solution, then $\Sigma_{+}$ is Lipschitz.
Proof.
For any $x\in\pi_{1}(\mathcal{NS}(u))$, there exists $y\in\mathbb{R}$ such that $(x,y)\in\mathcal{NS}(u)$. Since $u$ is a weak K.A.M. solution, we have $(x,y)\in\mathcal{NS}(T^{-}[u])$. By (a) of Lemma 2.12, we know $u$ is differentiable at $x$ and $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)=-\partial_{1}S(x,y)$. Since $S$ is ferromagnetic, we have $y=\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x))$, where $F$ is the lift of $f$ associated to $S$. On the other hand, By $(x,y)\in\mathcal{NS}(u)$ and Remark 2.11, we have $y=\Sigma_{+}(x)$. Therefore, we have
$\Sigma_{+}(x)=\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x)),\quad\forall x\in\pi_{1}(\mathcal{NS}(u)).$
Then we claim that $\Sigma_{+}$ is also Lipschitz on $\pi_{1}(\mathcal{NS}(u))$. In fact, for any $x_{1},x_{2}\in\pi_{1}(\mathcal{NS}(u))$,
$|\Sigma_{+}(x_{1})-\Sigma_{+}(x_{2})|$ $=\left|\pi_{1}\circ F(x_{1},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1}))-\pi_{1}\circ F(x_{2},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2}))\right|$
$\leq K^{\prime}\|(x_{1}-x_{2},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(x_{2}))\|_{\mathbb{R}^{2}}$
$\leq K^{\prime}\sqrt{1+K^{2}}|x_{1}-x_{2}|.$
where
$K^{\prime}=\sup_{x\in[0,1],\ |p|\leq K}\|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}(\pi_{1}\circ F)(x,p)\|_{(\mathbb{R}^{2})^{*}}.$
Now we show that $\Sigma_{+}$ is Lipschitz on whole $\mathbb{R}$. For any $x_{1},x_{2}\in\mathbb{R}$, by Corollary 3, there exist $x_{1}^{\prime\prime},x_{2}^{\prime}\in\pi_{1}(\mathcal{NS}(u))$ such that
$x_{1}\leq x_{1}^{\prime\prime}\leq x_{2}^{\prime}\leq x_{2},\quad\Sigma_{+}(x_{1}^{\prime\prime})=\Sigma_{+}(x_{1})\quad\text{ and }\quad\Sigma_{+}(x_{2}^{\prime})=\Sigma_{+}(x_{2}).$
Then
$|\Sigma_{+}(x_{1})-\Sigma_{+}(x_{2})|$ $=|\Sigma_{+}(x_{1}^{\prime\prime})-\Sigma_{+}(x_{2}^{\prime})|$
$\leq K^{\prime}\sqrt{1+K^{2}}|x_{1}^{\prime\prime}-x_{2}^{\prime}|$
$\leq K^{\prime}\sqrt{1+K^{2}}|x_{1}-x_{2}|.$
Hence $\Sigma_{+}$ is Lipschitz. ∎
Proposition 3.10.
Let $u$ be a weak K.A.M. solution. If $F$ is the lift of $f$ associated to $S$, then
$\forall x\in\mathbb{R},\quad\Sigma_{+}(x)=\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)).$
Proof.
Take $x\in\mathbb{R}$ and let $y=\Sigma_{+}(x)$. Then $T^{+}[u](x)=u(y)-S(x,y)$. Using (b) of Lemma 2.5, we have that
$-\partial_{1}S(x,y)\in\nabla^{-}T^{+}[u](x).$
Since $T^{+}[u]$ is differentiable, by (iv) if Proposition 2.2, we know that $\nabla^{-}T^{+}[u](x)=\{\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)\}.$ Hence we get
$-\partial_{1}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x).$
Since $S$ is ferromagnetic, if $F$ is the lift of $f$ associated to $S$, then
$y=\pi_{1}\circ F(x,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)).$
The proof ends. ∎
Corollary 4.
If $u$ is a discrete weak K.A.M. solution, then $T^{+}[u]$ is a $C^{1,1}$ function.
Proof.
We only need to show that $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u]$ is Lipschitz. Using Proposition 3.10 and the ferromagnetic condition of $S$, we have
$\ \mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x)=-\partial_{1}S(x,\Sigma_{+}(x)).$
Then for any $x_{1},x_{2}\in\mathbb{R}$,
$\left|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x_{1})-\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x_{2})\right|$ $=\left|\partial_{1}S(x_{1},\Sigma_{+}(x_{1}))-\partial_{1}S(x_{2},\Sigma_{+}(x_{2}))\right|$
$\leq K^{\prime\prime}\|(x_{1}-x_{2},\Sigma_{+}(x_{1})-\Sigma_{+}(x_{2}))\|_{\mathbb{R}^{2}}$
$\leq K^{\prime\prime}\sqrt{1+K^{\prime 2}+(K^{\prime}K)^{2}}|x_{1}-x_{2}|,$
where
$K^{\prime\prime}=\sup_{x\in[0,1],\ |y|\leq\|\Sigma_{+}\|_{C^{0}[0,1]}}\|\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}(\partial_{1}S)(x,y)\|_{(\mathbb{R}^{2})^{*}}.$
Hence $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u]$ is Lipschitz. ∎
3.3. $\Sigma_{+}$ propagates singularities
Proposition 3.11.
Let $u$ be a weak K.A.M. solution. Then $y\in\operatorname{Sing}(u)$ if and only if $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed bounded interval.
Proof.
Since $u$ is a weak K.A.M. solution, we know $u=T^{-}u-\bar{S}$. If $y\in\operatorname{Sing}(u)$, then $T^{-}[u]$ is also non-differentiable at $y$. Using (a) of Lemma 2.6, we have that there exist $x_{1}\neq x_{2}$ such that $T^{-}[u](y)=u(x_{i})+S(x_{i},y),\ \forall i\in\{1,2\}$. Using that $u$ is a weak K.A.M. solution again, we claim
$u(y)+\bar{S}=T^{-}[u](y)=u(x_{i})+S(x_{i},y),\ \forall i\in\{1,2\}.$
That is, $(x_{i},y)\in\mathcal{NS}(u),\ \forall i\in\{1,2\}.$ Hence $T^{+}[u](x_{i})=u(y)-S(x_{i},y)$ by Remark 2.11. That is, $\{x_{1},x_{2}\}\subset(\Sigma_{+})^{-1}(y)$. By Proposition 3.6, $(\Sigma)^{-1}(y)$ must be a non-degenerate closed bounded interval.
If $(\Sigma)^{-1}(y)$ is a non-degenerate closed bounded interval $[x_{-},x_{+}]$, then by Proposition 3.6, we know that $(x_{-},y),(x_{+},y)\in\mathcal{NS}(u)$. Hence $T^{-}[u](y)=u(x_{\pm})+S(x_{\pm},y)$ by Remark 2.11. By (a) of Lemma 2.6, we have $T-[u]$ is non-differentiable at $y$. Hence $y\in\operatorname{Sing}(u)$. ∎
Proposition 3.12.
Let $u$ be a weak K.A.M. solution. If $x\in\operatorname{Sing}(u)$, then $\Sigma_{+}(x)\in\operatorname{Sing}(u)$.
Proof.
Let $y=\Sigma_{+}(x)$. By Corollary 3, there exists $x^{\prime}$ such that $\Sigma_{+}(x^{\prime})=y$ and $(x^{\prime},y)\in\mathcal{NS}(u)$. Since $u$ is a weak K.A.M. solution, we know $u=T^{-}u-\bar{S}$. Hence $(x^{\prime},y)\in\mathcal{NS}(T^{-}[u])$ and it follows from (a) of Lemma 2.12 that $u$ is differentiable at $x^{\prime}$. Then $x\neq x^{\prime}$ since $x\in\operatorname{Sing}(u)$. Hence we claim that $(\Sigma)^{-1}(y)$ is not a singleton because it contains at least $x^{\prime}$ and $x$ By Proposition 3.6, $(\Sigma)^{-1}(y)$ must be a non-degenerate closed bounded interval. By Proposition 3.11, we have $y\in\operatorname{Sing}(u)$. ∎
3.4. $\Sigma_{+}$ admits a rotation number
The following lemma is classical. The reader may refer to any textbook on rotation theory. For the sake of self-containedness, we provide the proof here. We do not make any claim of originality for this lemma.
Lemma 3.13.
Let Φ:normal-:normal-$\Phi:\mathbb{R}\rightarrow\mathbb{R}$ be a continuous map such that
@@ 3.4. $\Sigma_{+}$ admits a rotation number
Using the property (i), by induction, one can show
$\forall x\in\mathbb{R},m\in\mathbb{N}_{+},\quad\Phi^{m}(x+1)=\Phi^{m}(x)+1.$
Hence $\Phi^{m}-\operatorname{id}$ is $1$-periodic and by continuity of $\Phi$ we can denote
@@
Using the property (ii), we claim that
$0\leq\beta_{m}-\alpha_{m}\leq 1$
In fact, for any $x_{1}\leq x_{2}\leq x_{1}+1$, we have that $\Phi^{m}(x_{1})\leq\Phi^{m}(x_{2})\leq\Phi^{m}(x_{1})+1$. Then
$\Phi^{m}(x_{1})-(x_{1}+1)\leq\Phi^{m}(x_{2})-x_{2}\leq\Phi^{m}(x_{1})+1-x_{1}.$
Hence $|(\Phi^{m}(x_{1})-x_{1})-(\Phi^{m}(x_{2})-x_{2})|\leq 1$ and thus $\beta_{m}-\alpha_{m}\leq 1$.
For any $n\in\mathbb{N}_{+},n\geq m$, we know that
$n=qm+r,\quad 0\leq r<m.$
It follows from $\forall y\in\mathbb{R},\ \alpha_{m}\leq\Phi^{m}(y)-y\leq\beta_{m},$ that
$\forall p\in\mathbb{N}_{+},\quad\alpha_{m}\leq\Phi^{pm}(y)-\Phi^{(p-1)m}(y)\leq\beta_{m}.$
Adding these inequalities for $p=1,2,\dots,q$, we get
$q\alpha_{m}\leq\Phi^{qm}(y)-y\leq q\beta_{m}.$
In particular, we have that
$\begin{cases}q\alpha_{m}\leq\Phi^{qm+r}(x)-\Phi^{r}(x)\leq q\beta_{m};\\ r\alpha_{1}\leq\Phi^{r}(x)-x\leq r\beta_{1}.\end{cases}$
Adding these two inequalities together and dividing them by $n$, we get
$\frac{q\alpha_{m}+r\alpha_{1}}{n}\leq\frac{1}{n}\left(\Phi^{n}(x)-x\right)\leq\frac{q\beta_{m}+r\beta_{1}}{n}.$
Since
$\left|\frac{q\beta_{m}+r\beta_{1}}{n}-\frac{q\alpha_{m}+r\alpha_{1}}{n}\right|$ $\rightarrow\frac{\beta_{m}-\alpha_{m}}{m}\quad(n\to+\infty),$
$\frac{\beta_{m}-\alpha_{m}}{m}$ $\rightarrow 0\quad\quad\quad\quad\ \ (m\to+\infty),$
@@
By combining Section 3.1 with Lemma 3.13, we proved that $\Sigma_{+}$ admits a rotation number $\rho(\Sigma_{+})$ for positive iterations. Moreover, for any $(x,p)\in\mathcal{A}^{*}$, by Proposition 3.10 we have
$\Sigma_{+}(x)=\pi_{1}\circ F(x,p).$
Hence $\rho(\Sigma_{+})$ equals to the rotation number of the Aubry set.
4. An explanation of Arnaud’s observation
In this section, we aim to prove Theorem 1.3. We first provide a formula of $(\Sigma_{+})^{-1}$.
@@ 4. An explanation of Arnaud’s observation
Let $u$ be a discrete weak K.A.M. solution. If $F$ is the lift of $f$ associated to $S$, then
$\forall y\in\mathbb{R},\quad(\Sigma_{+})^{-1}(y)=\pi_{1}\circ F^{-1}(y,\nabla^{+}u(y)).$
@@
Case 1: $u$ is differentiable at $y$. By Proposition 3.11, we know that $(\Sigma_{+})^{-1}(y)=\{x\}$ is a singleton. Using (b) of Lemma 2.5, we claim $\partial_{2}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)$. Hence the definition of generating function implies $x=\pi_{1}\circ F^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)).$
Case 2: $u$ is non-differentiable at $y$. By Proposition 3.11, we know that $(\Sigma_{+})^{-1}(y)$ is a non-degenerate closed interval $[x_{-},x_{+}]$. According to the proof of Proposition 3.6, we can find an increasing sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $x_{n}\rightarrow x_{-}$ and $u$ is differentiable at $y_{n}:=\Sigma_{+}(x_{n})$ for all $n\in\mathbb{Z}_{+}$. Then
$x_{n}=\pi_{1}\circ F^{-1}(y_{n},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y_{n}))$
by Case 1 and we claim that

### v2

@@ 3.3. $\Sigma_{+}[u]$ propagates singularities
Suppose that $y\in\operatorname{Sing}(u)$. Since $u=T^{-}[v]$ is non-differentiable at $y$, by part (a) of Lemma 2.22, there exist distinct $x_{1}\neq x_{2}$ such that $\{x_{1},x_{2}\}\subset\Sigma_{-}[v](y)$. Using Lemma 2.5, it follows that $\{x_{1},x_{2}\}\subset(\Sigma_{+}[u])^{-1}(y)$. By Lemma 3.4, $(\Sigma_{+}[u])^{-1}(y)$ must be a non-degenerate closed bounded interval.
Suppose that $(\Sigma_{+}[u])^{-1}(y)$ is a non-degenerate closed bounded interval $[x_{-},x_{+}]$. By Lemma 3.4, $x_{-},x_{+}\in\Sigma_{-}[v](y)$. Using part (a) of Lemma 2.22, $T^{-}[v]$ is non-differentiable at $y$. Hence, $y\in\operatorname{Sing}(u)$. ∎
Proposition 3.7.
Let $u=T^{-}[v]$, where $v\in SC(\mathbb{T})$. If $x\in\operatorname{Sing}(v)$, then $\Sigma_{+}[u](x)\in\operatorname{Sing}(u)$.
Proof.
Let $x\in\operatorname{Sing}(v)$ and $y=\Sigma_{+}[u](x)$. Take any $x^{\prime}\in\Sigma_{-}[v](y)$. By Lemma 2.21, $v$ is differentiable at $x^{\prime}$, so $x^{\prime}\neq x$. By Lemma 2.5, $\Sigma_{+}[u](x^{\prime})=y$, implying $\{x,x^{\prime}\}\subset(\Sigma_{+}[u])^{-1}(y)$. By Lemma 3.4, $(\Sigma_{+}[u])^{-1}(y)$ is a non-degenerate closed bounded interval. Hence by Proposition 3.6, $y\in\operatorname{Sing}(u)$. ∎
Corollary 3.8.
Let $u$ be a discrete weak K.A.M. solution. If $x\in\operatorname{Sing}(u)$, then $\Sigma_{+}[u](x)\in\operatorname{Sing}(u)$.
3.4. $\Sigma_{+}[u]$ admits a rotation number
The following lemma is well-known and we provide its proof for selfcontainedness.
@@ 3.4. $\Sigma_{+}[u]$ admits a rotation number
Let $\Phi:\mathbb{R}\rightarrow\mathbb{R}$ be a continuous map such that
@@
From property (i), by induction, we have:
$\Phi^{m}(x+1)=\Phi^{m}(x)+1,\quad\forall x\in\mathbb{R},m\in\mathbb{N}_{+}.$
Thus, $\Phi^{m}-\operatorname{id}$ is $1$-periodic. Define
@@
From property (ii), for any $x_{1}\leq x_{2}\leq x_{1}+1$, we have:
$\Phi^{m}(x_{1})\leq\Phi^{m}(x_{2})\leq\Phi^{m}(x_{1})+1.$
This implies
$|(\Phi^{m}(x_{1})-x_{1})-(\Phi^{m}(x_{2})-x_{2})|\leq 1,$
and hence $\beta_{m}-\alpha_{m}\leq 1$.
Let $n=qm+r$, where $0\leq r<m$. Then, using $\alpha_{m}\leq\Phi^{m}(x)-x\leq\beta_{m},$ we obtain:
$q\alpha_{m}\leq\Phi^{qm}(x)-x\leq q\beta_{m}.$
Additionally,
$r\alpha_{1}\leq\Phi^{r}(x)-x\leq r\beta_{1}.$
Combining these, we have:
$q\alpha_{m}+r\alpha_{1}\leq\Phi^{n}(x)-x\leq q\beta_{m}+r\beta_{1}.$
Dividing through by $n$, we get:
$\frac{q\alpha_{m}+r\alpha_{1}}{n}\leq\frac{\Phi^{n}(x)-x}{n}\leq\frac{q\beta_{m}+r\beta_{1}}{n}.$
@@
By combining Proposition 3.3 and Lemma 3.9, we proved that if $u\in T^{-}(SC(\mathbb{T}))$, then $\Sigma_{+}[u]$ admits a rotation number $\rho(\Sigma_{+}[u])$ for positive iterations for any $u\in T^{-}[SC(\mathbb{T})]$.
@@
If $u\in T^{-}(SC(\mathbb{T}))$ and $u$ is a sub-action, then $\rho(\Sigma_{+}[u])$ is the rotation number of the Aubry set.
@@
Let $\{x_{n}\}_{n\in\mathbb{Z}}$ be a $u$-calibrated sequence contained in the projected Aubry set. For any $n\in\mathbb{Z}$, we have:
$u(x_{n+1})-S(x_{n},x_{n+1})=u(x_{n})-\overline{S}.$
Since $u$ is a sub-action, for any $y\in\mathbb{R}$:
$u(x_{n})-\overline{S}\geq u(y)-S(x_{n},y).$
Thus:
$u(x_{n+1})-S(x_{n},x_{n+1})\geq u(y)-S(x_{n},y),$
which implies $x_{n+1}\in\Sigma_{+}[u](x_{n})$. By Lemma 3.2, $x_{n+1}=\Sigma_{+}[u](x_{n})$. The rotation number is given by:
$\lim_{n\to+\infty}\frac{1}{n}(\Sigma_{+}^{n}[u](x_{0})-x_{0})=\lim_{n\to+\infty}\frac{1}{n}(x_{n}-x_{0}).$
This is the rotation number of the Aubry set. ∎
Corollary 3.11.
Let $u$ be a discrete weak K.A.M. solution. Then $\operatorname{Sing}(u)$ is a forward invariant set under $\Sigma_{+}[u]$, with the same rotation number as the Aubry set.
4. Proofs of Theorems 1.4-1.7
In this section, we study several consequences of the dynamics $\Sigma_{+}[u]$ and prove the associated theorems. We operate under the assumption that the function $S$ satisfies Hypothesis 1, Hypothesis 2, and Hypothesis 3. The structure of this section is as follows:
  • In Section 4.1, we prove Theorem 1.4.
  • In Section 4.2, we prove Theorem 1.5.
  • In Section 4.3, we prove Theorems 1.6 and 1.7.
By analyzing the structure and properties related to the dynamics $\Sigma_{+}[u]$, we aim to uncover its significant behaviors and applications under different scenarios.
4.1. Regularization results
This result serves as a discrete analogue of Patrick Bernard’s regularization theorem [Ber07]. Notice that [Zav23] states that if $u$ is a discrete weak K.A.M. solution, then $T^{+}[u]$ is $C^{1}$.
Proposition 4.1.
If $v\in SC(\mathbb{T})$, then $T^{+}\circ T^{-}[v]$ is $C^{1}$. Moreover, if the discrete standard map $F$ is locally bi-Lipschitz, then $T^{+}\circ T^{-}[v]$ is $C^{1,1}$.
Proof.
From Lemma 2.22 (item (b)), for any $x\in\mathbb{R}^{d}$, $\Sigma_{+}[u](x)$ is a singleton if and only if $T^{+}[u]$ is differentiable at $x$. Let $u:=T^{-}[v]$. By Lemmas 3.2 and 2.22, $\Sigma_{+}[u](x)$ is a singleton for all $x\in\mathbb{R}^{d}$. This implies that $T^{+}[u]$ is differentiable everywhere. From item (b) of Lemma 2.19, we have:
(5) $-\partial_{1}S(x,\Sigma_{+}[u](x))=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u](x),\quad\forall x\in\mathbb{R}.$
Since $\partial_{1}S$ and $\Sigma_{+}[u]$ are continuous, it follows that $T^{+}[u]$ is $C^{1}$.
Assume $F$ is locally bi-Lipschitz, meaning both $F$ and $F^{-1}$ are locally Lipschitz continuous. Recall that
$\partial_{1}S(x,y)=-\pi_{2}\circ F^{-1}(y,\pi_{2}\circ F(x,y)).$
Since $F^{-1}$ is locally Lipschitz continuous, $\partial_{1}S$ is also locally Lipschitz continuous.
Denote $M:=\|\Sigma_{+}[u]\|_{C^{0}[0,1]}$. On the compact set $[0,1]\times[-M,M]$, $\partial_{1}S$ is Lipschitz continuous. From Lemma 3.5, $\Sigma_{+}[u]$ is Lipschitz continuous on $[0,1]$. Using the equation (5), we conclude that $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u]$ is Lipschitz continuous on $[0,1]$. Since $T^{+}[u]$ is $1$-periodic, the Lipschitz continuity of $\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}T^{+}[u]$ extends to $\mathbb{R}$. Thus, $T^{+}[u]$ is $C^{1,1}$ globally. ∎
Corollary 4.2.
If the discrete standard map $F$ is locally bi-Lipschitz, then for any bounded function $u:\mathbb{T}\rightarrow\mathbb{R}$, $T^{+}\circ T^{-2}[u]$ is $C^{1,1}$.
4.2. An explanation of Arnaud’s observation
To prove Theorem 1.5, we first derive an explicit formula for $(\Sigma_{+}[u])^{-1}$ as stated in Lemma 4.3.
Lemma 4.3.
Let $u=T^{-}[v]$ with $v\in SC(\mathbb{T})$ and assume $F$ is the discrete standard map. Then for any $y\in\mathbb{R}$,
$(\Sigma_{+}[u])^{-1}(y)=\pi_{1}\circ F^{-1}(y,\nabla^{+}u(y)).$
Proof.
We consider two cases based on the differentiability of $u$ at $y$.
Case 1: $u$ is differentiable at $y$. By Proposition 3.6, $(\Sigma_{+}[u])^{-1}(y)=\{x\}$, a singleton. From Lemma 2.19 (b), $\partial_{2}S(x,y)=\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)$. Using the definition of the discrete standard map (see Remark 1.9), we conclude:
$x=\pi_{1}\circ F^{-1}(y,\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y)).$
Case 2: $u$ is non-differentiable at $y$. By Proposition 3.6, $(\Sigma_{+}[u])^{-1}(y)$ is a non-degenerate closed interval $[x_{-},x_{+}]$. From the proof of Lemma 3.4, we can construct an increasing sequence $\{x_{n}\}_{n\in\mathbb{Z}_{+}}$ such that $x_{n}\rightarrow x_{-}$ and $u$ is differentiable at $y_{n}:=\Sigma_{+}[u](x_{n})$. By Case 1:
$x_{n}=\pi_{1}\circ F^{-1}(y_{n},\mathop{}\hphantom{\mskip-3.0mu}\mathrm{d}u(y_{n})).$
Taking limits, we obtain